A trader needs ₦300,000 to improve her business, she deposits ₦110,000 in her savings account at 5% per annum compound interest. she then adds ₦50,000 to her savings at the end of each year.
a) Find her total savings after 3 years.
b) By how much is her savings greater or lesser than ₦300,000 at the end of the third year.

Answer by Theo(13342)

she deposits 110,000.
at end of first year, she has 110,000 * 1.05 = 115,500.
she adds 50,000 to get 165,500.
at end of second year, she has 165,500 * 1.05 = 173,775.
she adds 50,000 to get 223,775.
at end of third year, she has 223,775 * 1.05 - 234963.75.
she adds 50,000 to get 284,963.75.
she needs 300,000
she is short by 300,000 - 284,963.75 = 15,036.25